OpenStudy (anonymous):

What would the derivative be for 1/(4x^2)?

zepdrix (zepdrix):

\[\Large \left(\frac{1}{4x^2}\right)'\qquad=\qquad \frac{1}{4}\left(x^{-2}\right)'\]Understand how to apply the Power Rule from here? :)

OpenStudy (anonymous):

Sweet!! Thanks! Jip, understand from here. :)